Physics307L F09:People/Rivera/Notebook/Speed Of Light
Speed of Light Lab Summary
Start with cardboard tube with photomultiplier in one end and LED emitter in the other. Connect DC power supply (150V-200V) to the LED emitter. Connect BNC cable from the LED emitter to the start in the TAC. Connect DC power supply (1800V-2000V) to the photo multiplier. Connect photo multiplier to the time delay input. Connect from photo multiplier output to stop on the TAC. Connect TAC output to oscilloscope.
We spent the first day setting up the experiment and testing the individual parts to make sure they were working. We were unable to get any data that day but did get some ideas and things to think about for the next week.
Before the next Monday we looked at Anne's Notebook and got some ideas from what they did on Wednesday.
On Monday we began taking data but noted a problem with the readings we were getting from the scope. We tried switching out the time delay but were still getting the anomalies. Next instead of using a delay we added more BNC cables to create a delay. After this we got better numbers but they were still fluctuating quite a bit.
We averaged the oscilloscope and took average readings using the peak output from the photomultiplier (this is achieved by rotating the multiplier to get the right polarization of the light) and averaging the reading from the scope. There were a lot of problems with this method as it was dependent on our ability to correctly turn the multiplier and taking a visual average from the scope. We also tried splitting the output from the multiplier into channel 2 on the scope but still got inconsistent numbers with what was expected.
Given these numbers we determined the speed of light to be **2x10^8 m/s (+/- 1x10^8)** as the mean of our readings.
Notes and Charts
Problems and Issuessee comment
The first problem we had was the setup. We spent a lot of time trying to locate all the parts need for the experiment. This ended up taking most of the first day.
Next it seemed that some of the equipment wasn't working correctly and had to change some of it out and just live with some of it.
It also seemed that we were getting some feedback from the LED signal that seemed to be causing the fluctuations in our data. Even with terminators the stream was quite erratic.In averaging the oscilloscope we add in another discrepancy to our data not really knowing how accurate the average is at only 128 cycle averaging. If we could of gotten a reading every tick in a file, and averaged the data ourselves on a spread sheet, I would have felt better about the numbers.see comment